OpenStudy (anonymous):

I am currently learning cofunction identities have to find tan(Alpha-Beta). Tan Alpha is 4/3 and tan beta is 15/8. I already have the equation set up. But I'm stuck at the denominator of the fraction. On the numerator, I have (4/3)-(15/8) On the denominator, It is 1+ (4/3)(15/8). I am stuck on how to solve this out because I am no good on fractions lol.

OpenStudy (atlas):

Solve the num and den separately

OpenStudy (atlas):

num (take lcm and simplify) = 4/3 - 15/8 = (8X4)- (15X3)/24 = (32-45)/24 = -13/24

OpenStudy (atlas):

similarly solve for denominator and divide

OpenStudy (anonymous):

I've simplified it already. just the 1 + 60/24 part on the denominator is what stumps me. Do i multiply 24 to 1 so I can give it a fraction